# This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible

### v3

@@ V Limitation
  • Hyperparameter Optimization: The Machine Learning models analyzed in this study were evaluated using their default hyperparameters, as outlined in the methodology. A comprehensive hyperparameter tuning process for each model–scaler–dataset combination was beyond the current scope; however, such optimization could potentially uncover different optimal pairings or further improve model performance.
@@
VI Conclusion
This comprehensive empirical study investigated the impact of 12 feature scaling techniques on 14 Machine Learning algorithms in 16 classification and regression datasets. Key findings reaffirmed the robustness of ensemble methods (e.g., Random Forest, gradient boosting family), which, along with Naive Bayes, largely maintained high performance irrespective of scaling. This offers efficiency gains by potentially avoiding preprocessing overhead. In stark contrast, models such as Logistic Regression, SVMs, MLPs, K-Nearest Neighbor, and TabNet demonstrated high sensitivity, with their performance critically dependent on scaler choice — a pattern consistent across both task types. Computational analysis also indicated that scaling choices can influence training/inference times and memory usage, with certain scalers being notably more resource-intensive.
This study contributes with one of the first systematic evaluations of such an extensive array of models, some less common models like TabNet, and scaling techniques — including transformations, e.g Tanh (TT) and Hyperbolic Tangent (HT), which are less commonly benchmarked as general-purpose scalers —, all within a unified Python framework. This is particularly relevant given that feature scaling is often applied in the literature without clear rationale, sometimes incorrectly before data splitting, which can lead to data leakage, or without verifying algorithm-specific benefits. By providing broad empirical evidence, our work offers clear guidance on how to mitigate these common issues, promoting informed scaling selection and more rigorous experimental design.
@@ VI Conclusion
Appendix A Tables Results
TABLE XII: Accuracy by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
@@ Appendix A Tables Results
TABLE XII: Accuracy by dataset, model, and scaling method.
TABLE XIII: Time to Train (s) by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
@@
TABLE XIII: Time to Train (s) by dataset, model, and scaling method.
TABLE XIV: Time to Inference (s) by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
@@
TABLE XIV: Time to Inference (s) by dataset, model, and scaling method.
TABLE XV: $R^{2}$ Score by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
@@
TABLE XV: $R^{2}$ Score by dataset, model, and scaling method.
TABLE XVI: MSE by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
@@
TABLE XVI: MSE by dataset, model, and scaling method.
TABLE XVII: MAE by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
@@
TABLE XVII: MAE by dataset, model, and scaling method.
TABLE XVIII: Time to Train (s) by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
@@
TABLE XVIII: Time to Train (s) by dataset, model, and scaling method.
TABLE XIX: Time to Inference (s) by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
@@
TABLE XIX: Time to Inference (s) by dataset, model, and scaling method.
TABLE XX: Memory Usage (kB) per Dataset and Scaling Method
Dataset NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
@@
TABLE XX: Memory Usage (kB) per Dataset and Scaling Method

### v4

@@ VI Limitation
  • Hyperparameter Optimization: The Machine Learning models analyzed in this study were evaluated using their default hyperparameters, as outlined in the methodology. A comprehensive hyperparameter tuning process for each model, scaler and dataset combination was beyond the current scope; however, such optimization could potentially uncover different optimal pairings or further improve model performance.
@@
  • K-Fold Cross-Validation: Although cross-validation is a widely recommended practice to obtain more reliable performance estimates, in this study, it was not employed due to the large number of datasets, models, and scaling configurations evaluated. Performing k-fold cross-validation in this experimental setup would substantially increase the computational cost, making the analysis impractical within the available resources. Instead, we adopted a hold-out validation strategy, which allowed for a consistent and feasible comparison across all scenarios. Future work could incorporate cross-validation or repeated runs to further validate the robustness of the findings.
@@
VII Conclusion
This comprehensive empirical study investigated the impact of 12 feature scaling techniques on 14 Machine Learning algorithms in 16 classification and regression datasets. Key findings reaffirmed the robustness of ensemble methods (e.g., Random Forest, gradient boosting family), which, along with Naive Bayes, largely maintained high performance irrespective of scaling. This offers efficiency gains by potentially avoiding preprocessing overhead. In stark contrast, models such as Logistic Regression, SVMs, MLPs, K-Nearest Neighbor, and TabNet demonstrated high sensitivity, with their performance critically dependent on scaler choice. Computational analysis also indicated that scaling choices can influence training/inference times and memory usage, with certain scalers being notably more resource-intensive.
This study contributes with one of the first systematic evaluations of such an extensive array of models, some less common models like TabNet, and scaling techniques — including transformations, e.g Tanh (TT) and Hyperbolic Tangent (HT), which are less commonly benchmarked as general-purpose scalers —, all within a unified Python framework. This is particularly relevant given that feature scaling is often applied in the literature without a clear rationale, sometimes incorrectly before data splitting, which can lead to data leakage, or without verifying algorithm-specific benefits. By providing broad empirical evidence, our work offers clear guidance on how to mitigate these common issues, promoting informed scaling selection and more rigorous experimental design.
@@ VII Conclusion
Appendix A Tables & Figures Results
Figure 4: $R^{2}$ results for the 7 datasets. The ensemble models are: RF, LGBM, CatBoost, and XGBoost.
Figure 5: MAE results for the 7 datasets. The ensemble models are: RF, LGBM, CatBoost, and XGBoost.
Figure 6: Time to train for the classification tasks. The ensemble models are: RF, LGBM, CatBoost, and XGBoost.
Figure 7: Time to inference for the classification tasks. The ensemble models are: RF, LGBM, CatBoost, and XGBoost.
Figure 8: Time to train for the regression tasks. The ensemble models are: RF, LGBM, CatBoost, and XGBoost.
Figure 9: Time to inference for the regression tasks. The ensemble models are: RF, LGBM, CatBoost, and XGBoost.
Figure 10: Memory Usage (kB) per Dataset and Scaling Method
TABLE V: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for accuracy.
Model Wilcoxon stat Wilcoxon $p$ Wilcoxon sig. Friedman stat Friedman $p$ Friedman sig.
LR 1927.5000 0.3228 No 44.9579 0.0000 Yes
SVM 1392.0000 0.0001 Yes 42.0466 0.0000 Yes
MLP 744.5000 0.0000 Yes 34.4050 0.0006 Yes
RF 368.0000 0.9710 No 10.2824 0.5912 No
NB 706.0000 0.3379 No 5.7230 0.9294 No
CART 164.0000 0.0372 No 5.5087 0.9388 No
LGBM 336.0000 0.0252 No 17.3499 0.1369 No
Ada 0.0000 0.3173 No 12.0000 0.4457 No
CatBoost 31.0000 0.0913 No 9.7680 0.6363 No
XGBoost 19.0000 0.0329 No 5.9406 0.9190 No
KNN 372.0000 0.0000 Yes 28.9849 0.0040 Yes
TabNet 2312.5000 0.0732 No 17.1439 0.1443 No
TABLE V: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for accuracy.
TABLE VI: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for $R^{2}$.
Model Wilcoxon stat Wilcoxon $p$ Wilcoxon sig. Friedman stat Friedman $p$ Friedman sig.
SVR 574.0000 0.0000 Yes 33.9245 0.0007 Yes
MLP 1193.0000 0.0083 Yes 47.3783 0.0000 Yes
RF 1721.0000 0.7753 No 13.2075 0.3541 No
CART 1152.0000 0.0530 No 5.9805 0.9171 No
LGBM 740.0000 0.0127 No 0.4284 1.0000 No
Ada 732.0000 0.0000 Yes 18.9670 0.0893 No
CatBoost 1463.0000 0.1510 No 2.5755 0.9979 No
LinearRegression 896.0000 0.1806 No 13.7553 0.3166 No
XGBoost 4.0000 0.7150 No 18.0000 0.1157 No
KNN 1275.0000 0.1860 No 21.8466 0.0393 No
TabNet 537.0000 0.0000 Yes 17.1805 0.1429 No
TABLE VI: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for $R^{2}$.
TABLE VII: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for MSE.
Model Wilcoxon stat Wilcoxon $p$ Wilcoxon sig. Friedman stat Friedman $p$ Friedman sig.
SVR 826.0000 0.0000 Yes 33.9245 0.0007 Yes
MLP 1187.0000 0.0077 Yes 47.3783 0.0000 Yes
RF 1650.0000 0.5471 No 13.2075 0.3541 No
CART 1209.0000 0.0987 No 5.9805 0.9171 No
LGBM 875.0000 0.0991 No 0.4284 1.0000 No
Ada 794.0000 0.0000 Yes 18.9670 0.0893 No
CatBoost 1347.0000 0.0508 No 2.5755 0.9979 No
LinearRegression 1273.5000 0.4236 No 13.7403 0.3176 No
XGBoost 4.0000 0.7150 No 18.0000 0.1157 No
KNN 1458.0000 0.6811 No 21.8466 0.0393 No
TabNet 811.0000 0.0000 Yes 17.1805 0.1429 No
TABLE VII: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for MSE.
TABLE VIII: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for MAE.
Model Wilcoxon stat Wilcoxon $p$ Wilcoxon sig. Friedman stat Friedman $p$ Friedman sig.
SVR 885.0000 0.0001 Yes 19.3019 0.0815 No
MLP 1382.0000 0.0723 No 32.5526 0.0011 Yes
RF 1234.0000 0.0140 No 6.9434 0.8613 No
CART 1217.5000 0.2036 No 5.5396 0.9375 No
LGBM 1138.0000 0.9950 No 5.3525 0.9452 No
Ada 1171.0000 0.0212 No 20.7147 0.0547 No
CatBoost 1255.0000 0.0181 No 9.9273 0.6223 No
LinearRegression 1237.5000 0.6677 No 6.0974 0.9111 No
XGBoost 4.0000 0.7150 No 18.0000 0.1157 No
KNN 1091.0000 0.0252 No 11.2601 0.5068 No
TabNet 697.0000 0.0000 Yes 13.6954 0.3206 No
TABLE VIII: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for MAE.
TABLE IX: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for time to train.
Model Wilcoxon stat Wilcoxon $p$ Wilcoxon sig. Friedman stat Friedman $p$ Friedman sig.
LR 648.5000 0.0000 Yes 39.2741 0.0001 Yes
SVM 1424.5000 0.0000 Yes 49.5541 0.0000 Yes
MLP 4831.5000 0.0000 Yes 44.8613 0.0000 Yes
RF 6612.5000 0.0006 Yes 14.2508 0.2850 No
NB 2654.0000 0.9371 No 21.3755 0.0451 No
CART 9137.0000 0.9677 No 21.1547 0.0482 No
LGBM 8306.5000 0.2143 No 15.7088 0.2049 No
Ada 6118.0000 0.0000 Yes 14.1374 0.2920 No
CatBoost 8183.0000 0.1609 No 8.5467 0.7411 No
XGBoost 4454.0000 0.0000 Yes 37.8845 0.0002 Yes
KNN 5243.5000 0.0000 Yes 84.2967 0.0000 Yes
TabNet 7290.0000 0.0105 No 21.1319 0.0485 No
SVR 494.0000 0.0000 Yes 45.6641 0.0000 Yes
LinearRegression 211.0000 0.0000 Yes 42.2801 0.0000 Yes
TABLE IX: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for time to train.
TABLE X: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for time to inference.
Model Wilcoxon stat Wilcoxon $p$ Wilcoxon sig. Friedman stat Friedman $p$ Friedman sig.
LR 648.5000 0.0000 Yes 39.2741 0.0001 Yes
SVM 1424.5000 0.0000 Yes 49.5541 0.0000 Yes
MLP 4831.5000 0.0000 Yes 44.8613 0.0000 Yes
RF 6612.5000 0.0006 Yes 14.2508 0.2850 No
NB 2654.0000 0.9371 No 21.3755 0.0451 No
CART 9137.0000 0.9677 No 21.1547 0.0482 No
LGBM 8306.5000 0.2143 No 15.7088 0.2049 No
Ada 6118.0000 0.0000 Yes 14.1374 0.2920 No
CatBoost 8183.0000 0.1609 No 8.5467 0.7411 No
XGBoost 4454.0000 0.0000 Yes 37.8845 0.0002 Yes
KNN 5243.5000 0.0000 Yes 84.2967 0.0000 Yes
TabNet 7290.0000 0.0105 No 21.1319 0.0485 No
SVR 494.0000 0.0000 Yes 45.6641 0.0000 Yes
LinearRegression 211.0000 0.0000 Yes 42.2801 0.0000 Yes
TABLE X: Wilcoxon test results (scaled vs. unscaled features) and Friedman test results (across all scaling methods) for time to inference.
TABLE XI: Accuracy by dataset, model, and scaling method.
Dataset Model NO $M$ MA ZSN PS VAST MC RS QT DS $T$ LS HT
@@ Appendix A Tables & Figures Results
TABLE XI: Accuracy by dataset, model, and scaling method.
TABLE XII: Time to Train (s) by dataset, model, and scaling method.
Dataset Model NO $M$ MA ZSN PS VAST MC RS QT DS $T$ LS HT
@@
TABLE XII: Time to Train (s) by dataset, model, and scaling method.
TABLE XIII: Time to Inference (s) by dataset, model, and scaling method.
Dataset Model NO $M$ MA ZSN PS VAST MC RS QT DS $T$ LS HT
@@
TABLE XIII: Time to Inference (s) by dataset, model, and scaling method.
TABLE XIV: $R^{2}$ Score by dataset, model, and scaling method.
Dataset Model NO $M$ MA ZSN PS VAST MC RS QT DS $T$ LS HT
@@
TABLE XIV: $R^{2}$ Score by dataset, model, and scaling method.
TABLE XV: MSE by dataset, model, and scaling method.
Dataset Model NO $M$ MA ZSN PS VAST MC RS QT DS $T$ LS HT
@@
TABLE XV: MSE by dataset, model, and scaling method.
TABLE XVI: MAE by dataset, model, and scaling method.
Dataset Model NO $M$ MA ZSN PS VAST MC RS QT DS $T$ LS HT
@@
TABLE XVI: MAE by dataset, model, and scaling method.
TABLE XVII: Time to Train (s) by dataset, model, and scaling method.
Dataset Model NO $M$ MA ZSN PS VAST MC RS QT DS $T$ LS HT
@@
TABLE XVII: Time to Train (s) by dataset, model, and scaling method.
TABLE XVIII: Time to Inference (s) by dataset, model, and scaling method.
Dataset Model NO $M$ MA ZSN PS VAST MC RS QT DS $T$ LS HT
@@
TABLE XVIII: Time to Inference (s) by dataset, model, and scaling method.
TABLE XIX: Memory Usage (kB) per Dataset and Scaling Method
Dataset NO $M$ MA ZSN PS VAST MC RS QT DS $T$ LS HT
@@
TABLE XIX: Memory Usage (kB) per Dataset and Scaling Method